Passing through the point (5,10) and parallel to the line with equation 7x+5y=11
Answer by stanbon(75887)

Passing through the point (5,10) and parallel to the line with equation 7x+5y=11
-------------
Find the slope of the given line:
y = (-7/5)x + (11/5)
slope = -7/5
---------------------
Equation Form: y = mx+b
Substitute and solve for "b":
10 =(-7/5)(5)+b
10 = -7 + b
b = 17
--------
Equation:
y = (-7/5)x + 17
=================
Cheers,
Stan H.